OpenStudy (anonymous):

Which expression matches the written rule below, where n is the term number? Each term in a sequence is equal to the term number increased by thirty-seven. 37 - n 37n n + 37 n - 37

OpenStudy (lucaz):

each term is equal to the term number: a_1=1, a_2=2, a_3=3... a_n=n increased by 37; a_1=1+37, a_2=2+37, a_3=3+37... a_n=n+37

OpenStudy (anonymous):

so wait ts n+37?

OpenStudy (anonymous):

thanks @lucaz

OpenStudy (lucaz):

yes.. no problem